# NavM[Uncaptioned image]rph: A Self-Evolving World Model for Vision-and-Language Navigation in Continuous Environments

### v1

@@ Abstract
Vision-and-Language Navigation in Continuous Environments (VLN-CE) requires agents to execute sequential navigation actions in complex environments guided by natural language instructions. Current approaches often struggle with generalizing to novel environments and adapting to ongoing changes during navigation. Inspired by human cognition, we present NavMorph, a self-evolving world model framework that enhances environmental understanding and decision-making in VLN-CE tasks. NavMorph employs compact latent representations to model environmental dynamics, equipping agents with foresight for adaptive planning and policy refinement. By integrating a novel Contextual Evolution Memory, NavMorph leverages scene-contextual information to support effective navigation while maintaining online adaptability. Extensive experiments demonstrate that our method achieves notable performance improvements on popular VLN-CE benchmarks. Code is available at this https URL.
@@ 4 Experiments
We evaluate our model on R2R-CE and RxR-CE [41] datasets, which transform discrete paths of R2R [5] and RxR [43] datasets into continuous environments. Experiments and ablation studies show our effectiveness. Please refer to Supplementary Material for source code, detailed experimental settings, more results and qualitative analysis.
@@ 4.3 Further Remarks
Other Ablation Studies. Refer to Supplementary Material for other experiments, such as detailed ablation study of self-evolution and CEM, comparison with other online testing strategies, hyper-parameter analysis (e.g., $T_{p}$), etc.

### v2

@@ Abstract
Vision-and-Language Navigation in Continuous Environments (VLN-CE) requires agents to execute sequential navigation actions in complex environments guided by natural language instructions. Current approaches often struggle with generalizing to novel environments and adapting to ongoing changes during navigation. Inspired by human cognition, we present NavMorph, a self-evolving world model framework that enhances environmental understanding and decision-making in VLN-CE. NavMorph employs compact latent representations to model environmental dynamics, equipping agents with foresight for adaptive planning and policy refinement. By integrating a novel Contextual Evolution Memory, NavMorph leverages scene-contextual information to support effective navigation while maintaining online adaptability. Extensive experiments demonstrate that NavMorph achieves notable performance improvements on popular VLN-CE benchmarks. Our Code is available at https://github.com/Feliciaxyao/NavMorph.
@@ 4 Experiments
We evaluate our model on R2R-CE and RxR-CE [41] datasets, which transform discrete paths of R2R [5] and RxR [43] datasets into continuous environments. Experiments and ablation studies show our effectiveness.
@@ 5 Conclusion
Acknowledgment
This work was supported in part by the National Key Research and Development Plan of China under Grant 2023YFC3310700, in part by the National Natural Science Foundation of China under Grants 62036012, U21B2044, 62236008, 62472422, and U2333215, and in part by Beijing Natural Science Foundation under Grant 4242051.
@@ 9.3 Other Ablation Studies
Computational Analysis. Table 4 in main paper demonstrates that NavMorph maintains comparable inference efficiency to the base model, with an average episode time of 21.22s vs 20.53s, while achieving nearly 4% improvements in both SR and SPL. In terms of parameter overhead, our CEM introduces only a marginal increase—adding 2.30M parameters compared to the base model’s 228.96M, accounting for merely 1.0% of the total model size. Given the consistent performance improvements, the computational and memory overheads are lightweight and acceptable.
@@
9.4 Extensive Discussion
While our method shares similarities with test-time adaptation (TTA), NavMorph fundamentally extends beyond this paradigm. Traditional TTA typically applies gradient-based parameter updates during inference for static classification tasks. In contrast, our approach incorporates a contextual evolution mechanism (CEM) within the RSSM framework, explicitly modeling environment state transitions during both training and inference phases. This mechanism enables the agent to adapt proactively to dynamic environments—not just during inference—by selectively integrating new scene observations while retaining historically relevant information.
At the core of this design lies the Contextual Evolution Memory (CEM), which enhances long-term reasoning by dynamically maintaining latent scene representations. Rather than accumulating all past experiences, CEM performs top-K scene retrieval based on visual similarity, maintaining memory entries that are most pertinent to the current context. This suppresses noisy or outdated trajectories and enhances the agent’s ability to infer plausible future transitions, particularly in out-of-distribution or evolving scenes, as often encountered in VLN-CE tasks. We provide ablations (Table 8) comparing variants of NavMorph with and without online evolution (NavMorph w/o SE*), confirming that the model remains effective even without runtime adaptation (SR: Base Model 43.77 → NavMorph w/o SE* 46.87 → NavMorph 47.91).
Importantly, our contribution lies not in the memory design itself, but in showing that adaptive evolution—when tightly integrated into a world model—can effectively improve navigation performance in VLN-CE literature.